1/4x-5/2=3/4x+3/2
Can you please show steps to sovling

Answer by stanbon(75887)

1/4x-5/2=3/4x+3/2
---
Multiply thru by 4 to get:
x - 10 = 3x + 6
2x = -16
x = -8
========
Cheers,
Stan H.